Solve for maturity value, discount period, bank discount, and proceeds. Assume a bank discount rate of 9%. Use the ordinary interest method. (Use Days in a year table.) (Do not round intermediate calculations. Round your final answers to the nearest cent.) 
Answer #1

Answer:

Face Value = 26300

bank discount rate or rate of interest = 9%

Length of note = 65 days

So, interest = 26300* 9% * 65/360 = $427.38

Maturity Value = 26300+ 427.38= $26727.38

Discount Period = 34-65 = 31 days

Bank Discount = $26727.38 * 9% * 31/360 = $207.14

Proceeds = Maturity Value - Bank Discount = $26727.38 - $207.14= $26934.52
Answer #2

Face Value P = $26300

Bank discount rate D =9%

Rate of interest R = 9%

Length of note LN = 65 days

Since Interest I =PRT

Therefore Interest I = $26300(9/100)(65/360) = $427.375=$427.38

Since Maturity Value MV = P+I 

Therefore Maturity Value MV = $26300+ $427.38= $26727.38

Since Discount Period DP = LN - (Day of discount in table - day of note in days table) 

Therefore  Discount Period DP = 65 - (110-76) = 65-34 = 31 days

Since, Bank Discount DB = (MV)(D)((DP)

Therefore Bank Discount DB = $26727.38(9/100)(31/360) = $207.137195 = $207.14

Since Proceeds PS = MV - BD 

Therefore Proceeds PS = $26727.38 - $207.14= $26520.24